OpenStudy (anonymous):

can someone explain to me Implicit Differentiation: if y^3 = 25x^2, determine dx/dt when x = 5 and dy/dt = 1

OpenStudy (anonymous):

these are the steps but i don't understand them :/ 3y^2*dy/dt = 50x*dx/dt When x = 5, y^3 = 25*5^2 y^3 = 625 y = cube root 625 = 8.55 3(8.55)^3*1 = 50*5*dx/dt 1875 = 250dx/dt 7.5 = dx/dt

OpenStudy (anonymous):

isn't y=2.92402?

OpenStudy (anonymous):

i did this on the calculator: sqrt (625)^(1/2)

OpenStudy (anonymous):

(1/3)**

OpenStudy (anonymous):

@satellite73

OpenStudy (agent0smith):

Let's start with this part... \[\Large \frac{ d }{ dt } y^3\]use the chain rule to differentiate it... so bring down the exponent, reduce the exponent by one, then multiply by the derivative of y w. respect to t \[\Large 3 y^2 *\frac{ dy }{ dt }\]

OpenStudy (anonymous):

so I can't replace the dy/dt by 1 since it's given?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Well yes but i was seeing if you understood the differentiation... don't worry about plugging in numbers till later.

OpenStudy (anonymous):

and 3y^2 is already the derivative, why do we need to use the chain rule?

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Because you have to multiply by the derivative of y with respect to t.

OpenStudy (agent0smith):

3y^2 is NOT the derivative of y^3, UNLESS you're just differentiating with respect to y.

OpenStudy (anonymous):

normally we have y dependent on x here y, and x both are dependent on some t, so we dont know what the function is so instead of y^3 >>> 3y^2 we have the chain rule y^3 >>> 3y^2 * y'

OpenStudy (agent0smith):

It's the same with dy/dx... the derivative of y with respect to x is not just 1... it's 1*dy/dx

OpenStudy (anonymous):

ok... :/ it's a little confusing..

OpenStudy (ranga):

agent0smith explained it nicely: \[\frac{ d }{ dy }(y ^{3}) =3y ^{2}\] But \[\frac{ d }{ dt }y ^{3} = \frac{ d }{ dy }y ^{3}\frac{ dy }{ dt } = 3y ^{2}\frac{ dy }{ dt }\]

OpenStudy (agent0smith):

It's the same thing with the chain rule \[\huge \frac{ d }{ dx } f(x)^n = n*f(x)^{n-1} * f \prime (x)\] you have to always multiply by the derivative at the end.

OpenStudy (anonymous):

why can't we plug in the numbers right away if it's given?

OpenStudy (agent0smith):

You can... you're probably just better off not doing it until you really know what you're doing.

OpenStudy (anonymous):

right i understand that... @ranga

OpenStudy (anonymous):

well i was taught to plug in.. that's why.

OpenStudy (anonymous):

that's the confusing part.

OpenStudy (agent0smith):

Fair enough... but you can't plug in until you have this step: 3y^2*dy/dt = 50x*dx/dt

OpenStudy (anonymous):

exactly what i have right now.

OpenStudy (anonymous):

so dy/dt is 1...

OpenStudy (anonymous):

in the steps above, why is this: 3(8.55)^3*1 = 50*5*dx/dt?

OpenStudy (anonymous):

y=8.55? y is suppose to be 2.92402.

OpenStudy (anonymous):

and it's raised to the 3rd, and why not 2?

OpenStudy (agent0smith):

When x = 5, y^3 = 25*5^2 y^3 = 625 y = cube root 625 = 8.55 i think the next line after that is a mistake and it should by to the power of 2

OpenStudy (anonymous):

so y=2.92402 is correct?

OpenStudy (agent0smith):

y^3 = 625 y = cube root 625 = 8.55

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

i did this in the calculator: sort(625)^(1/3)

OpenStudy (anonymous):

sqrt*

OpenStudy (agent0smith):

And you'll get 8.55... 2.94 cubed is close to 3^3 which is 27. Not 625.

OpenStudy (anonymous):

i swear im getting 2.92402

OpenStudy (anonymous):

i plugged it in exactly like iwrote it.

OpenStudy (agent0smith):

sort(625)^(1/3) why are you taking the square root of a cubed root...

OpenStudy (agent0smith):

y^3 = 625 y = 625^(1/3)

OpenStudy (anonymous):

OHHH

OpenStudy (anonymous):

i thought the sqrt could do cube root.

OpenStudy (anonymous):

if we did the exponent also.

OpenStudy (agent0smith):

That would mean ((625^(1/3))^(1/2) = 625^(1/6)

OpenStudy (anonymous):

thank you! i got it !